if three times the complement of an angle is subtracted by the supplement of the angle , then the result is ten . find the supplement of the complement
Answer by cleomenius(959)

180 - x - 3(90 - x) = 10
180 -x -270 + 3x = 10
-90 + 2x = 10
2x = 100
x = 50 degrees the angle
40 degrees is the complement
130 degrees is the supplement
3(40) = 130 - 10.
the result does check.
Cleomenius.